OpenStudy (anonymous):

The Polynomial P(s)=4s^5+s^3+7s^2+3 is a -What degree and -What roots?

OpenStudy (anonymous):

Well the degree is the highest power of s in this case its 5

OpenStudy (anonymous):

how are you supposed to find the roots? graph perhaps or newton's method?

OpenStudy (anonymous):

give me a sec and I will try to find the roots using synthetic division.

OpenStudy (anonymous):

checked using wolframalpha and there is only one irrational root so don't bother

OpenStudy (anonymous):

Oh ok there ya go :D

OpenStudy (anonymous):

not sure how to do it without graphing or something. you can check the possible rational zeros, which are \[\pm 1,\pm 3\pm \frac{1}{2}. \pm\frac{3}{2}, \pm\frac{1}{4},\pm\frac{3}{4}\]

OpenStudy (anonymous):

Ha WolframAlpha rocks :D

OpenStudy (anonymous):

but they don't work so don't bother!

OpenStudy (anonymous):

yeah especially for a fifth degree poly with 12 possible rational zeros. you feel like checking?

OpenStudy (anonymous):

Not any more, lol